# Jingxuan Li UCLA

### v2

@@ Appendix G Inducing VLM’s Persona Detailed Information
Figure 12: Two examples of trials evaluating VLM’s preferences. For each trial, a social media short video is used, and a screenshot is taken at the 2-second timestamp. The user then interacts with the VLM using a question-and-answer format to assess the model’s attitude toward the screenshot’s content. In these examples, TikTok content and the GPT4o model are used for demonstration.
@@
Figure 14: Example of Inducing the VLM’s Persona Using the ISQ Questionnaire Strategy: When the calculated score exceeds 60, the Vision-Language Model (VLM) chooses to stay engaged with the content for 45 seconds before scrolling down.
@@
Table 18: Value preference outcomes across different models and input settings on Value-Spectrum. The settings include direct multi-modal responses from Vision-Language Models (VLMs) and combinations of image descriptions generated by different VLMs with Large Language Models (LLMs).

### v3

@@ Appendix H Inducing VLM’s Persona Detailed Information
Value Yes (Image) Yes (Video) No (Image) No (Video) Not Sure (Image) Not Sure (Video)
Benevolence 82.67 76.00 16.00 16.67 1.33 7.33
Tradition 94.67 96.00 4.67 1.33 0.67 2.67
Self-Direction 95.33 92.67 2.67 2.67 2.00 4.67
Achievement 88.67 78.00 10.67 11.33 0.67 10.67
Conformity 82.67 80.00 14.67 11.33 2.67 8.67
Hedonism 87.33 83.33 12.67 10.67 0.00 6.00
Security 94.00 92.67 6.00 4.00 0.00 3.33
Stimulation 96.00 97.33 3.33 1.33 0.67 1.33
Power 87.33 82.67 11.33 7.33 1.33 10.00
Universalism 97.33 97.33 1.33 0.67 1.33 2.00
Average 90.60 87.60 8.40 8.30 1.00 5.58
Table 9: Human Annotation Results for Value Alignment of Screenshots and Videos (Percentages).
Figure 12: Two examples of trials evaluating VLM’s preferences. For each trial, a social media short video is used, and a screenshot is taken at the 2-second timestamp. The user then interacts with the VLM using a question-and-answer format to assess the model’s attitude toward the screenshot’s content. In these examples, TikTok content and the GPT-4o model are used for demonstration.
@@
Figure 14: Example of inducing the VLM’s persona using the ISQ strategy: When the calculated score exceeds 60, the Vision-Language Model (VLM) chooses to stay engaged with the content for 45 seconds before scrolling down.
@@
Table 10: Value preference outcomes across different models and input settings on Value-Spectrum. The settings include direct multi-modal responses from VLMs and combinations of image descriptions generated by different VLMs with LLMs.